If you obtain a loan of 800,000 pesos at the rate of 10% compounded annually in order to build a house, how much must you pay monthly to amortize the loan within a period of ten years starting at the end of first month?
Answer #1

10,382.04 pesos

Step-1:Calculation of equivalent monthly interest
(1+i)^n = (1+i)^n
(1+i)^12 = (1+0.10)^1
(1+i)^12 = 1.10
1+i = 1.10^ (1/12)
1+i = 1.00797414
i = 0.00797414
So,
Equivalent monthly interest rate = 0.00797414
Step-2:Calculation of monthly payment
Monthly Payment =-pmt(rate,nper,pv,fv)
= 10,382.04 pesos
Where,
rate = 0.00797414
nper = 10*12 = 120
pv = 8,00,000 pesos
fv = 0